OpenStudy (anonymous):

Find the equation (in slope-intercept form) of the line passing through the points with the given coordinates. (-5, -4), (-3, 4)

OpenStudy (anonymous):

m=(4-(-4))/(-3-(-5))=4 y=4x+c 4=4*(-3)+c c=16 equation is y=4x+16

OpenStudy (anonymous):

Where did the 16 come from? I got to 4x and got lost.

OpenStudy (anonymous):

after he got slope a more common way to find the equation is : y-y1=m(x-x1) where x1 and y1 are the coordinates for EITHER of the two points used to find the slope. so choosing the first: y-(-5)=4(x-(-4)) y+5=4x+16 y=4x+11

OpenStudy (anonymous):

except i messed up** should read y+4=4(x+5) y=4x+20-4 y=4x+16

OpenStudy (anonymous):

4=4*(-3)+c c=16